Dickinson Company has $12,020,000 million in assets. Currently half of these assets are financed with long- erm debt at 10.1 percent and half with common stock having a par value of $8. Ms. Smith, Vice President of Finance, wishes to analyze two refinancing plans, one with more debt (D) and one with more equity (E). The company eams a return on assets before interest and taxes of 10.1 percent. The tax rate is 40 percent. Tax loss carryover provisions apply, so negative tax amounts are permissable. Under Plan D, a $3,005,000 million long-term bond would be sold at an interest rate of 12.1 percent and 375,625 shares of stock would be purchased in the market at $8 per share and retired. Under Plan E, 375,625 shares of stock would be sold at $8 per share and the $3,005,000 in proceeds would be used to reduce long-term debt. a. How would each of these plans affect earnings per share? Consider the current plan and the two new plans. Round your answers to 2 decimal places.) Answer is complete and correct. Current Plan D Plan E Plan Eamings per share 0.48 0.39 0.48 us b-1. Compute the earnings per share if return on assets fell to 5.05 percent. (Negative amounts should be indicated by a minus sign. Round your answers to 2 decimal places.) Answer is complete but not entirely correct Current Plan D Plan E Eamings per share 2 3 4 5 6 9
b-3. Compute the earnings per share if return on assets increased to 15.1 percent. (Round your answers t decimal places.) 3Answer is complete but not entirely correct. Current PlanPlan DPlan E Eamings per share 13 21 0.96s 1.35 2.41 b-4. Which plan would be most favorable if return on assets increased to 15.1 percent? Consider the current plan and the two new plans Current Plan Plan E OOPlan D c-1. If the market price for common stock rose to $10 before the restructuring, compute the earnings per share. Continue to assume that $3,005,000 million in debt will be used to retire stock in Plan D and $3,005,000 million of new equity will be sold to retire debt in Plan E. Also assume that return on assets is 10.1 percent (Round your answers to 2 decimal places.) Answer is complete but not entirely correct. Current Plan D Plan E Earnings per share 0.61s 0.49$ 1.82 c-2. If the market price for common stock rose to $10 before the restructuring, which plan would then be most attractive? 2 3 4 5 9
Answer #1
a) Current Plan Plan D Plan E
EBIT (12020000*10.1%) 1214020 1214020 1214020
Interest:
= 6010000*10.1% = 607010
= 6010000*10.1%+3005000*12.1% = 970615
= (6010000-3005000)*10.1% = 303505
EBT 607010 243405 910515
Tax at 40% 242804 97362 364206
NI 364206 146043 546309
# of shares:
= 6010000/8 = 751250
= 751250-375625 = 375625
= 751250+375625 = 1126875
EPS $               0.48 $           0.39 $              0.48
b-1) Current Plan Plan D Plan E
EBIT (12020000*5.05%) 607010 607010 607010
Interest:
= 6010000*10.1% = 607010
= 6010000*10.1%+3005000*12.1% = 970615
= (6010000-3005000)*10.1% = 303505
EBT 0 -363605 303505
Tax at 40% 0 -145442 121402
NI 0 -218163 182103
# of shares:
= 6010000/8 = 751250
= 751250-375625 = 375625
= 751250+375625 = 1126875
EPS $                    -   $         -0.58 $              0.16
b-3) Current Plan Plan D Plan E
EBIT (12020000*15.10%) 1815020 1815020 1815020
Interest:
= 6010000*10.1% = 607010
= 6010000*10.1%+3005000*12.1% = 970615
= (6010000-3005000)*10.1% = 303505
EBT 1208010 844405 1511515
Tax at 40% 483204 337762 604606
NI 724806 506643 906909
# of shares:
= 6010000/8 = 751250
= 751250-375625 = 375625
= 751250+375625 = 1126875
EPS $               0.96 $           1.35 $              0.80
b-4) Plan D
c-1) Current Plan Plan D Plan E
EBIT (12020000*15.10%) 1815020 1815020 1815020
Interest:
= 6010000*10.1% = 607010
= 6010000*10.1%+3005000*12.1% = 970615
= (6010000-3005000)*10.1% = 303505
EBT 1208010 844405 1511515
Tax at 40% 483204 337762 604606
NI 724806 506643 906909
# of shares:
= 6010000/8 = 751250
= 751250-3005000/10 = 450750
= 751250+3005000/10 = 1051750
EPS $               0.96 $           1.12 $              0.86
c-2) Plan D